### v2

# A Narrative Review of Identity, Data, and Location Privacy Techniques in Edge Computing and Mobile Crowdsourcing

Abstract
As digital technology advances, the proliferation of connected devices poses significant challenges and opportunities in mobile crowdsourcing and edge computing. This narrative review focuses on the need for privacy protection in these fields, emphasizing the increasing importance of data security in a data-driven world. Through an analysis of contemporary academic literature, this review provides an understanding of the current trends and privacy concerns in mobile crowdsourcing and edge computing. We present insights and highlight advancements in privacy-preserving techniques, addressing identity, data, and location privacy. This survey serves as a useful resource for researchers, industry professionals, and policymakers, offering an overview of privacy challenges and potential solutions in these interconnected domains.
Keywords Privacy; Security; Edge Computing, Location, Mobile Crowd-sourcing, Cloud Computing
1 Introduction
The advent of the digital era has resulted in a significant surge in connected devices, generating a substantial volume of data transmitted over the Internet [1]. Managed predominantly via cloud infrastructures, the ubiquity of smart devices such as smartphones, tablets, smartwatches, and fitness trackers has woven them into the fabric of daily life. These devices gather vast amounts of contextual information about users, including their location, activities, and environmental conditions, crucial for applications designed to predict user behavior and provide personalized experiences [2, 3].
Mobile crowdsourcing has emerged as a significant phenomenon where individuals collectively contribute data through various digital platforms [4]. While applications such as traffic monitoring systems benefit from this crowdsourced data by offering real-time insights, the process raises substantial privacy concerns related to data transparency and the risks of unauthorized sensitive information access [5, 6].
Moreover, the exponential growth in mobile data traffic presents challenges in data management and processing [7]. Edge computing, which processes data closer to its source rather than relying solely on central servers, has been identified as a promising solution to alleviate network congestion. However, this decentralization also introduces potential privacy and security vulnerabilities that need to be addressed [8].
This paper explores the intersection of mobile crowdsourcing and edge computing with a particular focus on privacy preservation. We provide an overview of existing research addressing these technologies and their implications for privacy, aiming to equip researchers, practitioners, and policymakers with the necessary insights to navigate the complexities of ensuring privacy while advancing technological capabilities.
1.1 Literature Review Strategy
Our literature review strategy involved searches in academic repositories such as ACM Digital Library, IEEE Xplore, Springer, Elsevier, Google Scholar, DBLP, CiteSeerX, Microsoft Academic Search, Web of Science, and ScienceDirect. Publications from the period 2012 to 2021 were primarily considered, with a final selection of approximately 52 peer-reviewed papers from an initial pool of 150 papers based on relevance to mobile crowdsourcing and edge computing. These papers were analyzed for trends, methodologies, findings, and gaps in research.
Search Queries: Structured queries were employed to search for relevant literature, focusing on privacy preservation, data security, and the integration of mobile crowdsourcing with edge computing, among other aspects.
1. ("Mobile Crowdsourcing" AND "Edge Computing" AND "Privacy Preservation") AND ("Data Security" OR "Real-Time Data Processing") AND (2012-2021)
2. ("Mobile Crowdsourcing" AND "Edge Computing") AND ("User Context" OR "Personalized Services") AND ("IoT" OR "Cloud Computing") AND (peer-reviewed)
3. ("Mobile Crowdsourcing" AND "Edge Computing") AND ("Privacy" AND "Information Overload") AND ("Smart Devices" OR "User Behavior Prediction") AND (2012-2021)
1.2 Contributions
This work contributes to the field:
  1. 1.
    It offers an interdisciplinary examination of mobile crowdsourcing and edge computing, emphasizing the importance of privacy preservation.
  2. 2.
    It synthesizes both contemporary and foundational literature, providing an understanding of identity privacy, data privacy, and location privacy (based on their relevance to importance) in these interconnected domains, and identifying current research trends.
A meta-review of previous surveys with focus on privacy in these individual domains is presented in Table 1.
Table 1: Related Surveys.
Domain Focus Key Contributions and Privacy Aspects
Edge Computing Real-time and local data processing, IoT integration, Edge AI, Security Studies highlighted capabilities for local processing, integration with AI, and identified gaps in privacy protection specific to edge computing environments. Key studies include [9], [10], [11], [12], [13], [8], [14], [15].
Mobile Computing Security Trends, Healthcare, Cloud Integration, AR Applications Focused on evolving security challenges, particularly in mobile cloud contexts and healthcare applications during the COVID-19 pandemic, and AR enhancements. Relevant studies include [16], [17], [18], [19], [20].
1.3 Privacy Concerns in Mobile Crowdsourcing and Edge Computing
Next, we discuss mobile crowdsourcing and edge computing and examine the privacy challenges in these domains, focusing on identity privacy, data privacy, and location privacy due to their significant impact on protecting users in these interconnected systems.
Identity Privacy
Identity privacy refers to the protection of an individual’s identity from being exposed or inferred [21] during the process of mobile crowdsourcing or edge computing activities. This includes safeguarding personal information such as names, social security numbers, and other identifiers that can link activities to a specific individual.
Mobile crowdsourcing often involves collecting data from numerous participants. Ensuring that these participants remain anonymous or pseudonymous is crucial to prevent misuse of personal information [22]. Exposing personal identifiers can lead to identity theft, where malicious actors could use the stolen identities to commit fraud or other illegal activities [23]. Protecting identity privacy is fundamental in establishing and maintaining trust among users.
Data Privacy
Data privacy refers to the protection of the actual data collected, processed, and stored [24] during mobile crowdsourcing and edge computing. This involves ensuring that sensitive information remains confidential and is not accessed, altered, or disclosed without authorization.
Crowdsourced data can include sensitive information, such as health data, financial transactions, or personal communications. Protecting this data from unauthorized access and ensuring its integrity is highly important [24]. Various regulations, such as the General Data Protection Regulation (GDPR) or the Health Insurance Portability and Accountability Act (HIPAA), mandate strict data privacy measures. Data breaches can have severe consequences, including financial loss, reputational damage, and loss of user trust.
Location Privacy
Location privacy involves the protection of information regarding an individual’s geographical location obtained through mobile crowdsourcing and edge computing. This includes ensuring that location data cannot be used to track or monitor an individual’s movements without their consent.
Unauthorized access to location data can lead to stalking, harassment, or other malicious activities where an individual’s movements are tracked without their knowledge or consent [25]. Disclosure of sensitive locations, such as homes, workplaces, or places of interest, can pose security risks and privacy violations [26]. Users should have control over who can access their location data and for what purposes. It is, therefore, important to ensure the location privacy for the sake if users’ autonomy and their right to control their personal information [27].
2 Mobile Crowdsourcing
Mobile crowdsourcing represents a field where collective user contributions, primarily via mobile devices, drive data collection and service provision [28, 29]. By leveraging the ubiquitous nature of mobile devices, this domain taps into a vast pool of user-submitted data, encompassing geospatial and behavioral data as well as environmental and health metrics. Its applications are vast, spanning environmental monitoring, disaster response, urban planning, healthcare, and transportation, highlighting the versatility and impact of mobile crowdsourcing [30].
A quintessential example of the potential of mobile crowdsourcing is seen in Waze [31], a navigation app that integrates user-generated traffic data to provide real-time routing advice and traffic condition updates. This not only showcases the practical implementation of mobile crowdsourcing but also demonstrates how user participation can significantly enhance service quality and responsiveness.
Figure 1: Anatomy of a mobile crowdsourcing campaign
Key Stages in Mobile Crowdsourcing: We show the key stages in mobile crowdsourcing in Figure 1 .
  1. 1.
    Tasking Stage: This initial phase involves defining clear and actionable data collection objectives that align with the overarching goals of the crowdsourcing campaign.
  2. 2.
    Collecting Stage: At this juncture, data is actively acquired from users through their mobile devices. This data can range from environmental readings to user-generated content and location data, providing a rich dataset for analysis.
  3. 3.
    Storing Stage: The collected data is centralized on cloud or edge servers for processing. This stage is crucial for managing the volume of data and ensuring it is accessible for analysis.
  4. 4.
    Mining Stage: Here, advanced analytics and data mining techniques are applied to the aggregated data to extract meaningful insights, patterns, and trends.
  5. 5.
    Publishing Stage: The final stage involves disseminating the findings and insights to end-users or stakeholders, often through the same mobile platforms used for data collection.
Technological Frameworks and Challenges: The technological infrastructure supporting mobile crowdsourcing includes a range of components from mobile devices and network communication protocols to cloud and edge computing platforms. These technologies facilitate the efficient collection, transmission, and processing of crowdsourced data. However, this complex ecosystem also introduces challenges related to scalability, data quality control, and the integration of heterogeneous data sources.
Privacy Considerations: The expansive data collection inherent to mobile crowdsourcing initiatives raises significant privacy concerns. The transmission of sensitive personal information, such as location data and personal identifiers, to centralized servers poses risks related to data security and user privacy [32]. Addressing these concerns, the field is increasingly looking towards edge computing as a viable solution. By processing data closer to its source, edge computing offers enhanced performance, reduced transmission costs, and improved data security, presenting a promising avenue for mitigating privacy risks in mobile crowdsourcing [33].
Literature: In mobile computing, recent surveys have focused on security trends, healthcare applications during the COVID-19 pandemic, and the integration with cloud computing. A study [16] explores evolving security challenges in mobile cloud computing and strategies for safeguarding these environments. The significant role of mobile computing in healthcare, especially during the pandemic, is highlighted in a survey [17], addressing both challenges and potential solutions. Additionally, the integration of mobile computing with edge computing is examined in [18], where the authors discuss the benefits of leveraging edge computing for reducing latency and improving processing capabilities in mobile applications. The survey [19] delves into the advancements in mobile augmented reality (AR), focusing on the integration of mobile computing with AR to enhance user experiences and application performance. Further research [20] reviews the application of mobile computing in smart cities.
Challenges: As mobile crowdsourcing continues to evolve, future research and development efforts are likely to focus on enhancing privacy protections, improving data processing efficiencies, and expanding the applicability of crowdsourcing models. Innovations in encryption, decentralized data processing, and user-centric privacy controls are expected to play pivotal roles in shaping the next generation of mobile crowdsourcing solutions. Through the collective power of mobile device users, mobile crowdsourcing can bring transformative changes across numerous domains. However, ensuring the ethical and secure handling of user-contributed data remains a critical challenge that must be addressed to realize the full potential of this innovative approach.
2.1 Privacy Concerns in Mobile Crowdsourcing
2.1.1 Identity Privacy in Mobile Crowdsourcing
Identity privacy in mobile crowdsourcing is a crucial aspect of safeguarding user data, which includes sensitive information like names, addresses, and contact details. Discussions on identity privacy focus on the implementation of robust forward security measures to protect cloud data [34, 22]. Identity-based encryption methods are highlighted as a means to secure cloud data, showcasing the multifaceted approach required to enhance privacy protection in cloud computing contexts [23]. The Consolidated Identity Management Systems (CIDM) are proposed to safeguard user identity by separating authorization credentials, adding an extra authentication layer, and securing communication links with cloud service providers [35]. Further discussions emphasize blockchain-based trust models and strategies for identity management that negate the need for trusted third parties [36, 37].
2.2 Data Privacy in Mobile Crowdsourcing
Data privacy in mobile crowdsourcing involves developing methods and techniques focused on enhancing data security and privacy within cloud environments. Research efforts include sophisticated methods for conducting spatial data queries that prioritize user privacy without compromising data utility [38]. Techniques leveraging advanced cryptographic tools such as homomorphic encryption show the potential for employing encryption to fortify privacy measures in cloud computing [24].
2.3 Location Privacy in Mobile Crowdsourcing
Ensuring the privacy of location data in mobile cloud computing involves various innovative methods such as cipher-text retrieval techniques for location privacy protection [26]. Other approaches include the use of spatial cloaking to anonymize location data for queries in cloud environments [39]. Advanced cryptographic techniques, such as SHA-2 for robust mutual authentication and location privacy, are employed to ensure secure communication and privacy preservation in mobile cloud computing settings [40].
Collectively, these research efforts illustrate the diverse strategies being explored to protect identity, data, and location privacy in cloud computing. By leveraging encryption, spatial cloaking, and cryptographic authentication, these methods offer promising avenues for securing sensitive information against unauthorized access and misuse, thereby contributing to the broader goals of privacy and security in the digital age.
These methods are summarized in Table 2 .
Table 2: Privacy Concerns and Solutions in Mobile Cloud Computing
Privacy Concern Research Focus Key Solutions
Identity Privacy in Cloud Computing Safeguarding sensitive user data such as names, addresses, and contact details Robust forward security measures [34, 22] Identity-based encryption methods [23] Consolidated Identity Management Systems (CIDM) [35] Blockchain-based trust models and identity management [36, 37]
Data Privacy in Cloud Computing Enhancing data security and privacy within cloud environments Spatial data query methods prioritizing user privacy [38] Advanced cryptographic tools such as homomorphic encryption [24]
Location Privacy in Cloud Computing Ensuring privacy of location data Cipher-text retrieval techniques for location privacy protection [26] Spatial cloaking to anonymize location data [39] SHA-2 for robust mutual authentication and location privacy [40]
3 Edge Computing and Privacy Concerns
Edge computing represents a paradigm shift in data processing and management, emphasizing the decentralization of computational resources to bring data storage and processing closer to the sources of data [8]. This approach is particularly beneficial for real-time and latency-sensitive applications across diverse domains such as autonomous vehicles, smart homes, emergency response systems, smart agriculture, and modernized power grids.
Conceptual Framework: Edge computing seeks to address the inherent limitations of centralized cloud computing models by distributing computational tasks to the edge of the network [41]. This strategic distribution significantly reduces latency by minimizing the distance between the data source and the processing unit and alleviates bandwidth constraints on the network. Such enhancements boost overall system responsiveness and reliability.
Technological Advancements and Architectural Design: The architectural design of edge computing is inherently modular and scalable. As depicted in Figure 2, the typical architecture comprises three primary layers: the user layer, edge layer, and cloud layer. Each layer plays a distinct role in the ecosystem, with smart devices at the user layer generating data, edge servers at the edge layer processing this data in real-time, and cloud servers at the cloud layer providing long-term storage and further analytics capabilities.
Figure 2: Edge computing architecture illustrating the interplay between user, edge, and cloud layers.
Benefits of Edge Computing: The decentralization inherent in edge computing brings several benefits. By processing data closer to its source, it significantly reduces latency, which is crucial for applications that require real-time feedback, such as autonomous driving and emergency response systems [42]. Furthermore, edge computing contributes to bandwidth efficiency by minimizing the volume of data that needs to be transmitted to the cloud, thereby reducing network congestion and associated costs.
Integration with Smart Technologies: Edge computing is integral to the evolution of smart technologies [20]. It enables the seamless integration of IoT devices, smart sensors, and other data sources, facilitating the development of intelligent applications that can operate autonomously and interact with their environment in real-time. This integration is crucial for the advancement of smart cities, industrial automation, and personalized healthcare, among other applications.
Literature: Recent studies, such as those by [9] and [10], distinguish edge computing from cloud computing by detailing its capabilities for real-time and local data processing, highlighting emerging research, and practical applications. Research in [11] explores ’Edge AI’, examining how edge computing and AI are integrated, their applications, and the impact of AI algorithms on decision-making and data processing at the network edge. Issues of security and privacy specific to edge computing are addressed in [12], where it is noted that conventional cloud computing privacy methods are inadequate. [13] highlights concerns about location privacy in edge computing environments, while [8] reviews security and privacy preservation methods, noting a predominant focus on cloud computing over edge-specific strategies. Additionally, a survey [14] discusses edge computing’s influence on IoT development, advocating for its suitability over other computing paradigms for IoT scenarios. Furthermore, [15] presents a comprehensive review of current research on security in edge computing, pinpointing challenges in new models, application scenarios, and technological environments, with specific attention to access control, key management, privacy protection, attack mitigation, and anomaly detection.
3.1 Privacy Concerns in Edge Computing
Edge computing, despite its benefits, introduces complex privacy issues due to its decentralized nature. This subsection delves into identity, data, and location privacy, highlighting ongoing research and emerging solutions aimed at enhancing security and privacy in edge environments.
3.1.1 Identity Privacy in Edge Computing
Several studies have explored pseudonym technology, cryptographic methods, and data aggregation schemes in edge computing environments to preserve identity privacy [43, 44, 45]. However, these methods are challenged by de-anonymization attacks [46, 47, 48] and limited data integrity [44]. Notably, research efforts are being made to enhance security and privacy through various methods focused on maintaining anonymity, preventing unauthorized access, and ensuring user data integrity. Innovations include anonymous authentication mechanisms leveraging identity-based cryptography for secure device-to-edge server communications [49], and the combination of secret sharing with homomorphic encryption to bolster data privacy [50]. Another approach involves using fully homomorphic encryption with multikey capabilities to secure data processing [51], and optimizing federated learning to enhance privacy and efficiency in edge environments [52].
3.1.2 Data Privacy in Edge Computing
In edge computing, ensuring data privacy is crucial due to decentralized processing and storage. Innovative approaches include partitioning deep neural networks to balance privacy with energy consumption [53], and integrating federated learning with hybrid differential privacy and adaptive compression to protect data in industrial settings [54]. Challenges in location-based services are addressed by crafting specific privacy protection strategies [25], and a novel framework for secure data sharing emphasizes privacy-preserving computations across the computing landscape [55]. Additionally, Local Differential Privacy (LDP) techniques process data while preserving personal information, though they can reduce data utility due to noise addition [56].
3.1.3 Location Privacy in Edge Computing
Ensuring location privacy within edge computing is paramount due to the proliferation of location-based services. Techniques include task offloading with reinforcement learning to protect privacy in autonomous transport systems [41], and dual K-anonymous caching strategies within edge environments to safeguard location data [57]. Dynamic strategies to protect location privacy in multi-user mobile edge systems highlight the need for agile response mechanisms in privacy protection [58]. Additional efforts focus on balancing service efficiency with privacy in location-based services [25], and task offloading frameworks that consider location privacy [58]. Vehicular crowdsourcing applications also reflect a focus on driver privacy, combining pseudonym and mix-zone techniques [59, 60].
This literature review is summarized in Table 3.
Table 3: Privacy Concerns and Solutions in Edge Computing
Privacy Concern Research Focus Challenges and Solutions
Identity Privacy Pseudonym technology, cryptographic methods, data aggregation schemes [43, 44, 45] De-anonymization attacks [46, 47, 48] Limited data integrity [44] Anonymous authentication using identity-based cryptography [49] Secret sharing with homomorphic encryption [50] Fully homomorphic encryption with multikey capabilities [51] Optimized federated learning [52]
Data Privacy Ensuring data privacy due to decentralized processing and storage Partitioning deep neural networks for privacy and energy balance [53] Federated learning with hybrid differential privacy and adaptive compression [54] Location-based service privacy strategies [25] Secure data sharing frameworks [55] Local Differential Privacy (LDP) techniques [56]
Location Privacy Protecting privacy in location-based services Task offloading with reinforcement learning in autonomous transport systems [41] Dual K-anonymous caching strategies [57] Dynamic multi-user privacy protection [58] Balancing service efficiency with location privacy [25] Task offloading frameworks for location privacy [58] Pseudonym and mix-zone techniques in vehicular crowdsourcing [59, 60]
Challenges: Despite its numerous advantages, edge computing faces several challenges, including security concerns, data privacy issues, and management complexities [18]. The distributed nature of edge computing introduces multiple points of vulnerability that must be secured against potential threats. Managing data privacy also becomes more complex as data is processed across numerous edge devices and locations. Solutions to these challenges include implementing advanced encryption techniques, establishing robust access control measures, and developing standardized protocols for data management and security.
The increasing need to embed AI capabilities at the edge allows devices to perform sophisticated data analysis and decision-making processes locally. This enhances the autonomy of smart applications and reduces dependency on central servers. Furthermore, the ongoing development of 5G networks is expected to significantly bolster the capabilities of edge computing, enabling faster data processing and exchange rates.
4 Discussion
This paper has explored the critical privacy concerns associated with mobile crowdsourcing and edge computing. As these technologies continue to evolve and become integral to modern data processing and storage solutions, addressing privacy concerns becomes paramount.
Identity privacy in cloud and edge computing is a multifaceted challenge. Research emphasizes robust forward security measures and identity-based encryption to protect sensitive user information [34, 22, 23]. Proposals such as CIDM aim to enhance security by separating authorization credentials and securing communication links [35]. Additionally, blockchain-based trust models and decentralized identity management systems offer promising approaches to reduce the need for trusted third parties, mitigating the risks of identity theft and unauthorized access [36, 37]. However, significant challenges remain, particularly in countering de-anonymization attacks and ensuring the integrity of anonymized data [46]. Ongoing research into advanced cryptographic methods and innovative authentication mechanisms, including those leveraging identity-based cryptography, is needed for addressing these vulnerabilities.
Ensuring data privacy in cloud and edge computing environments is critical due to the decentralized nature of these technologies. Innovative approaches such as partitioning deep neural networks to balance privacy with energy consumption and employing federated learning with hybrid differential privacy are making significant strides in this area [53, 54]. Techniques like homomorphic encryption and LDP have shown potential in protecting data privacy by enabling secure data processing and analysis [24, 56]. Despite their promise, these techniques often introduce challenges such as increased computational overhead and reduced data utility due to noise addition.
The protection of location privacy is particularly crucial in mobile computing due to the pervasive use of location-based services. Research has proposed several methods to address this issue, including cipher-text retrieval techniques, spatial cloaking, and SHA-2 for robust mutual authentication [26, 39, 40]. Strategies such as task offloading with reinforcement learning and dual K-anonymous caching have been explored to enhance location privacy in edge computing environments [41, 57]. These techniques offer dynamic and scalable solutions to protect location data while maintaining the efficiency of edge services.
The implications of these privacy concerns are vast and multifaceted. As mobile computing becomes more ubiquitous, ensuring robust privacy protection mechanisms is not only a technical challenge but also a societal imperative. The interplay between user trust and technology adoption cannot be overlooked. Compliance with the regulations such as data protection laws requires the development of sophisticated privacy-preserving technologies that can adapt to diverse legal requirements and protect user data across different jurisdictions.
Future research should focus on enhancing the scalability and efficiency of privacy-preserving methods, addressing the challenges of de-anonymization and data utility, and exploring new cryptographic techniques to ensure robust privacy protection. The integration of artificial intelligence and machine learning into privacy protection strategies offers promising avenues for innovation. These technologies can help identify and mitigate privacy risks in real-time, providing an additional layer of security in dynamic and complex computing environments. Additionally, interdisciplinary collaboration among researchers, industry practitioners, and policymakers is essential to create standardized guidelines and frameworks for privacy protection in mobile computing.
5 Conclusion
This paper has examined the critical privacy concerns inherent in mobile crowdsourcing and edge computing, focusing on identity, data, and location privacy. Despite significant advancements, robust and scalable privacy-preserving methods are still required to address challenges such as de-anonymization attacks and data integrity issues. Ensuring user trust, complying with regulatory standards, and promoting the broader adoption of mobile computing technologies depend on effectively addressing these privacy concerns. Future research should prioritize developing scalable and efficient privacy-preserving techniques and integrating artificial intelligence and machine learning for real-time privacy risk mitigation. Additionally, interdisciplinary collaboration among researchers, industry practitioners, and policymakers is essential to create standardized guidelines and frameworks for privacy protection.